### v1

# Traffic Flow and Speed Monitoring Based On Optical Fiber Distributed Acoustic Sensor

Abstract
In the realm of intelligent transportation systems, accurate and reliable traffic monitoring is crucial. Traditional devices, such as cameras and lidars, face limitations in adverse weather conditions and complex traffic scenarios, prompting the need for more resilient technologies. This paper presents traffic flow monitoring method using optical fiber-based Distributed Acoustic Sensors (DAS). An innovative vehicle trajectory extraction algorithm is proposed to derive traffic flow statistics. In the processing of optical fiber waterfall diagrams, Butterworth low-pass filter and peaks location search method are employed to determine the entry position of vehicles. Subsequently, line-by-line matching algorithm is proposed to effectively track the trajectories. Experiments were conducted under both real highway and tunnel scenarios, showing that our approach not only extracts vehicle trajectories more accurately than the classical Hough and Radon transform-based methods, but also facilitates the calculation of traffic flow information using the low-cost acoustic sensors. It provides a new reliable means for traffic flow monitoring which can be integrated with existing methods.
Index Terms:
Optical Fiber Distributed Acoustic Sensor (DAS), Vehicle Detection, Trajectory Tracking, Vehicle Speed Estimation, Traffic Flow.
publicationid: pubid:
I Introduction
With the advancing opportunities in technology, intelligent transportation systems (ITS) have emerged as a promising direction for future transportation systems [1]. These systems can aid in monitoring and managing traffic flow using various tools such as cameras and sensors. Such technologies facilitate real-time traffic monitoring and control, reducing congestion and enhancing traffic flow [2]. Central to these are the components of vehicle detection and trajectory extraction. Vehicle detection involves recognizing and localizing target objects [3]. Vehicle trajectories, on the other hand, are derived from vehicle detection and tracking data gathered by devices on the road. Given the motion trajectory and environmental changes, accurately predicting a vehicle’s future trends and behaviors becomes vital for traffic monitoring [4]. Nonetheless, the efficient detection of features and precise prediction of a target vehicle’s trajectory remains challenging.
Figure 1: Comparison of camera and optical fiber observations. Cameras cannot deal well with these situations but fiber systems are not affected.
Nowadays, methods for monitoring traffic flow have predominantly turned to technologies such as cameras, radars, and lidars [5]. Deep Learning algorithms, including YOLO (2D multi-object detection) [6] and CenterPoint (3D multi-object detection) [7], are increasingly utilized for vehicle detection. Additionally, an effective combination of the Kalman filter and the Hungarian algorithm can be harnessed for state estimation and data association, thereby facilitating vehicle tracking and real-time positioning [8]. Nonetheless, as shown in Fig. 1.(a), the efficacy of these devices can be severely compromised by environmental conditions and their deployment strategies. Elements like rain, snow, fog, storms, and suboptimal lighting can distort or damage equipment, making round-the-clock, all-weather vehicle detection challenging. Compounding these issues, cameras, given their fixed viewing angles, introduce dead zones and issues relating to object sizes [9]. These problems are accentuated when detecting smaller objects nearby or larger objects at extended distances, as illustrated in Fig 1.(b). Present 2D detection methods often fall short in identifying distant small objects. While lidars can discern distance information over longer ranges, their high costs hinder widespread deployment, especially at highway ends [10]. Contrasting with cameras and lidars, which serve as “visual detectors”, underground optical fibers can be likened to “acoustic detectors”. Benefiting from numerous design advantages such as resistance to electromagnetic interference, heightened sensitivity, cost-effectiveness, durability, stealth, and adaptability, optical fibers may provide another reliable way for continuous and comprehensive traffic state perception in smart highways [11] and can be integrated with camera and Lidar-based methods.
In this study, we investigate methods using underground optical fiber-based Distributed Acoustic Sensing (DAS) for monitoring highway traffic, including highway and tunnel scenarios. The method mainly includes two parts: (1) vehicle detection and tracking by signal processing; and (2) traffic density, flow, and velocity estimation.
In the processing of optical fiber data, the raw data is presented as three-dimensional waterfall diagrams. Then image preprocessing through min-max normalization and wavelet threshold denoising is presented. Based on these preprocessed optical fiber waterfall diagrams, we propose a vehicle position detection algorithm based on the Butterworth low-pass filter and peaks location search. Then vehicle trajectories are extracted based on line-by-line matching and polynomial fitting, thus achieving real-time detection and overall tracking for vehicles. We then present the calculation methods for three traffic flow information (i.e. velocity, flow and density) on road profile and road segment respectively, realizing traffic flow and speed monitoring on highways.
The performances of the proposed methods are verified in real datasets collected from highway and tunnel scenarios. The results show that the vehicle tracking results extracted by our methods are more accurate than the traditional Hough and Radon transform-based methods. In the tunnel scenarios, our algorithm accuracy is up to 90.90%, which validates the usability of the proposed methods in practical scenarios considering the very low-cost signal processing based method. In complex vehicle conditions like congestion and illegal overtaking, the proposed methods can track vehicles when line overlapping and intersecting are present, validating the effectiveness of the proposed methods for complicated vehicle scenarios.
The main contributions of the paper are as follows:
  1. 1.
    We introduce the DAS system for vehicle detection, designed to gather data from both highways and tunnels.
  2. 2.
    We present a novel approach for vehicle detection and tracking, relying solely on optical fiber signals. This method exhibits remarkable accuracy, particularly within tunnel settings.
  3. 3.
    Our proposed technique effectively addresses challenges such as vehicle overtaking and congestion-induced crossover situations in tunnels.
II Related Works
Since distributed fiber optical sensing systems have been used in transportation [12], there have been some works using the DAS system for traffic monitoring, but these are mostly based on image and signal processing algorithms, such as Hough transform [13] and Radon transform [14]. Subsequent scholars are more committed to studying how to denoise the collected fiber signal, rather than analyzing the information in the fiber signal. In recent years, Deep Learning has made remarkable achievements in image denoising. An attention-based Convolutional Neural Network (CNN) structure has been introduced specifically for denoising short-time DAS signals [15]. However, these algorithms can only be used for denoising or line extracting, rather than monitoring changes in vehicle positions and real-time tracking. Additionally, their emphasis does not extend to the statistical analysis of traffic flow information. Moreover, DAS systems are commonly deployed in smart cities [16] and on campuses [17], where vehicles typically pass in sequence, and the situation is relatively straightforward. This configuration does not consider the challenges posed by long-distance installations on highways and the continuous tracking of vehicles over extended distances.
Most of the existing optical fiber signal processing relies on image processing techniques, such as the Hough transform and Radon transform, rather than processing at the pixel level. Because of the extensive radiation range of optical fiber signals, the Hough transform often identifies a single vehicle trajectory line as multiple lines. Conversely, the Radon transforms may overlook detecting lines with short driving distances. In addition, both the Hough transform and the Radon transform are primarily applicable for the detection of straight lines, making it challenging to discern instances of traffic congestion and speed variation.
In this paper, we propose a pixel-level approach to extract vehicle trajectories. Our designed method offers enhanced accuracy throughout the tracking process, effectively addressing challenges like crossovers and overlaps. Furthermore, we provide a comparative analysis of our proposed method with the Hough transform and Radon transform in Section V.
III Method
III-A Problem Formulation
The original optical fiber data recorded by the DAS system can be represented as a matrix denoted by A:
$A=\begin{pmatrix}a_{11}&a_{12}&\cdots&a_{1n}\\ a_{21}&a_{22}&\cdots&a_{2n}\\ \vdots&\vdots&\ddots&\vdots\\ a_{m1}&a_{m2}&\cdots&a_{mn}\end{pmatrix}=(a_{ij})$ (1)
where $m$ denotes the time points, $n$ denotes the distance points, and $a_{ij}\in\mathbb{R^{+}}$. To facilitate the visualization of vehicle data collected from optical fiber, we can draw optical fiber waterfall diagrams and display fiber stripes:
Fiber waterfall diagram: Three-dimensional waterfall diagrams after the visualization of the original optical fiber matrix A in Eq. (1) with the distance as the horizontal axis and time as the vertical axis.
Fiber stripe: Clear banded stripes visible to the human eye in a fiber waterfall diagram. In this paper, fiber stripes represent the driving trajectories of the vehicles.
According to the above definition, we can draw optical fiber waterfall diagrams based on the collected fiber signals. In this section, we provide a visual representation of optical fiber waterfall diagrams in highway tunnel datasets, as shown in Fig. 2. In Fig. 2, the brightly colored lines, denoted as fiber stripes in the optical fiber waterfall diagrams, signify the spatiotemporal positions of vehicles. Highways, particularly in tunnel scenarios, often present unique conditions, such as congestion(the merging lines depicted in Fig. 2 (a) and (b)) and overtaking(the crossing lines depicted in Fig. 2 (c)). Accurately distinguishing the trajectory of each vehicle and conducting traffic flow statistics is a challenge that requires exploration. This is the primary objective of this paper.
Figure 2: Optical fiber waterfall diagrams illustrating vehicle congestion in a 2000-meter section of a highway tunnel dataset: (a) five vehicles congested in a jam, (b) six vehicles congested in a jam, and (c) one car overtaking illegally.
III-B Signal Preprocessing
The optical fiber data distribution among different datasets may vary differently, so signal preprocessing is needed to eliminate the background noise as much as possible to obtain a cleaner signal. In this section, we employ two techniques for signal preprocessing: min-max normalization and weighted wavelet threshold denoising.
III-B1 Normalization
Min-max normalization is one of the most popular techniques of data normalization, which is a linear transformation of the original data, mapping the original value x into the value in the interval [0, 1]. The formula is as follows:
${}x_{\rm{norm}}=\frac{x-\rm{min}}{\rm{max}-\rm{min}}$ (2)
where $\rm{max}$ and $\rm{min}$ are the maximum and minimum values of the sample data respectively.
The standardization method in this paper is to process the whole optical fiber waterfall diagram, rather than a single row or single column; otherwise, it will affect the subsequent discrimination of trucks and cars, scaling the matrix corresponding to the whole diagram to [0,1].
III-B2 Signal Denoising
In signal processing, noise reduction is the primary task to be solved, and in our case, we employ a wavelet threshold denoising algorithm combining hard and soft threshold functions.
Due to its excellent time-frequency characteristics, wavelet theory has gained widespread use in signal processing and has experienced rapid development [18]. The energy of the signal corresponds to the wavelet coefficient with a larger amplitude, while the energy of noise corresponds to the wavelet coefficient with a smaller amplitude. Hence, an optimal threshold can be established, where wavelet coefficients exceeding this threshold are identified as valuable signals and are either retained or contracted. Conversely, wavelet coefficients falling below the threshold are deemed noise and are subsequently removed. Ultimately, the signal undergoes reconstruction through the inverse wavelet transform, resulting in the acquisition of a denoised signal.
For the selection of wavelet threshold function, a wavelet threshold denoising algorithm combining hard and soft threshold functions is employed to denoise vehicle vibration signals. The formula is as follows:
$\overline{{\omega}_{j,k}}=\begin{cases}{\omega}_{j,k}-a\lambda,&{\omega}_{j,k}\geq\lambda;\\ 0,&\lvert{\omega}_{j,k}\rvert\textless\lambda;\\ {\omega}_{j,k}+a\lambda,&{\omega}_{j,k}\leq-\lambda.\end{cases}$ (3)
where ${\omega}_{j,k}$ is the wavelet coefficient, and $\lambda$ is the threshold. When $a=0$, it is a hard threshold function; when $a=1$, it is a soft threshold function; when $0\textless a\textless 1$, it represents a compromise between the hard and soft thresholds. In this paper, according to the actual situation, an ideal denoised diagram can be obtained when $a=0.5$.
The normalized and denoised results of different wavelet threshold denoising algorithms are shown in Fig. 3. After denoising, both the peak value and fundamental trend of the vibration signal are effectively retained.
Figure 3: Comparisions among different wavelet threshold denoising algorithms. (a) normalized diagram, (b) a = 0, (c) a = 1, and (d) a = 0.5.
III-C Vehicle Position Detection
While the preprocessing steps can effectively eliminate global noise and interference, achieving accurate extraction of structured information, including vehicle positions from the optical fiber stripe, requires additional filtering. Therefore, we employ a real-time vehicle positioning approach through a peak location search algorithm.
III-C1 Smoothing method of Butterworth Filter
The Butterworth filter, named after the British engineer and physicist Stephen Butterworth [19], is characterized by a frequency response curve that is flat to the maximum in the passband (i.e., a bandpass filter) without a ripple, gradually rolling off to zero in the stopband. The Butterworth low-pass filter can be expressed by the following formula for amplitude squared relative to frequency:
$\lvert H(\omega)\rvert^{2}=\frac{1}{1+\epsilon^{2}(\frac{\omega}{\omega_{c}})^{2n}}$ (4)
where n is the order of the filter, $\omega_{c}$ is the cutoff frequency, and $\omega$ is the operating frequency (passband frequency) of circuit, $\epsilon$ is the maximum passband gain. The value of $\epsilon$ is determined by the following formula:
${}H_{1}=\frac{H_{0}}{\sqrt{1+\epsilon^{2}}}$ (5)
where $H_{0}$ and $H_{1}$ represent the maximum and minimum passband gains, respectively.
It’s worth noting that the higher the order of the Butterworth filter, the steeper the amplitude-frequency response, resulting in a more effective filtering effect.
III-C2 Peaks location search algorithm
According to the features of the wave peaks, satisfying the principle that the first derivative is zero and the second derivative is negative, the position of the wave peaks can be searched as follows:
(1) Suppose the projection curve can be expressed as $V=[v_{1},v_{2},\ldots,v_{n}]$.
(2) Calculate the first-order difference vector ${\rm{Diff}}_{V}$:
${\rm{Diff}}_{v}(i)=V(i+1)-V(i)$ (6)
where $i\in 1,2,\ldots,N-1$.
(3) Apply the sign function to the difference vector:
${\rm{Trend}}={\rm{sign}}({\rm{Diff}}_{v})$ (7)
where
${\rm{sign}}(x)=\begin{cases}1,&\text{if }x>0;\\ 0,&\text{if }x=0;\\ -1,&\text{if }x<0.\end{cases}$ (8)
(4) Iterate through the $\rm{Trend}$ vector from the tail to perform the following:
${\rm{Trend}}(i)=\begin{cases}1,\text{if }{\rm{Trend}}(i)=0\text{ and }{\rm{Trend}}(i+1)\geq 0;\\ -1,\text{if }{\rm{Trend}}(i)=0\text{ and }{\rm{Trend}}(i+1)<0.\end{cases}$ (9)
(5) Perform the first-order difference operation on the $\rm{Trend}$ vector, as in Step 2, to obtain $R=\rm{Diff}(\rm{Trend})$.
(6) Traverse the difference vector $R$. If $R(i)=-2$, then $i+1$ is a crest of the projection vector $V$, and the corresponding crest is $V(i+1)$. If $R(i)=2$, then $i+1$ is a trough of the projection vector $V$, and the corresponding trough is $V(i+1)$.
After applying the Butterworth low-pass filter to the waterfall figure column-wise and subsequently extracting the wave peaks, relatively accurate vehicle position points can be obtained. Fig. 4 displays the raw optical fiber signal within the first 5 minutes from the starting position of the optical fiber, along with the processing results. In the figure, the dummy line represents the original signal amplitude, the solid line represents the signal after the Butterworth filter, and the red points represent the detected peaks.
Figure 4: Raw fiber signal and filtered signal. (a) raw optical fiber signal of the starting position, (b) signal filtered by Butterworth Lowpass Filter and searched peaks.
The vehicle position detection algorithm flow is shown in Algorithm 1.
Input: $A=(a_{ij})$: original optical fiber matrix; $i=1,...,m$, $j=1,...,n$, $a_{ij}\in\mathbb{R^{+}}$: selected time and distance.
Output: $D=(d_{ij})$: preproccessed optical fiber matrix; $S$: Vehicle counts that can be extracted; $k_{s}\in\mathbb{N^{+}},s=1,...,S$: appearance time of each vehicle.
1 $B=(b_{ij})$leftarrow MinMaxNormalization($A=(a_{ij})$);
2 $D=(d_{ij})$leftarrow WaveletDenoising($B=(b_{ij})$);
3 $k_{s}$leftarrow FindPeak(ButterworthFilter($D_{1:m,1}$));
4 $(k_{s},1)$ is the first pair of time and distance position points of each extracted vehicle trajectory;
Return: Vehicle counts $S$; appearance time of each vehicle $k_{s}$; preprocessed optical fiber matrix $D=(d_{ij})$.
Algorithm 1 Position Detection Algorithm.
III-D Vehicle trajectory extraction
Figure 5: Workflow of the total algorithm.
In this section, we propose an algorithm for line-by-line matching based on the vehicle motion model. According to Newton’s first law, vehicles can typically be regarded as moving at a uniform velocity in the absence of lane changes, emergency stops, or overtaking maneuvers.
Considering the real highway scenario in this study, the space-time information reflected by optical fiber within the optical fiber waterfall diagram exhibits a straight-line pattern. This straight line’s slope is related to velocity, while the intercept is related to arrival time. This characteristic aligns well with a uniform linear motion model represented as:
$x=vt+x_{0}$ (10)
where $v$ is the vehicle’s velocity, $x_{0}$ is the initial position where the vehicle appears, and $x$ is the vehicle’s position at time $t$.
Figure 6: Trajectory extraction algorithm. The black line segment represents the velocity range and the red dot represents trajectory key points.
Therefore, after determining the starting position via vehicle position detection, knowing the vehicle’s velocity enables the determination of its position at different points along its trajectory. However, in the case of vehicle congestion, where each vehicle moves within a fixed range, there is a risk of multiple vehicles overlapping in their tracks. To address this challenge, we expand the speed range of each vehicle from a static line segment to a dynamic line segment. The selection of this dynamic line segment is determined by the slope of the line fitted using historical position and velocity, incorporating a certain degree of confidence. Fig. 6 illustrates the dynamic line segments within complex tunnel scenes, facilitating the tracking of key points in vehicle trajectories through line-by-line matching.
Given the key points of vehicle trajectories, we can reconstruct the complete vehicle trajectories and subsequently calculate the instantaneous speed of the vehicles. Furthermore, in scenarios where most short-distance vehicles on the highway maintain a consistent speed, we can further employ linear fitting to derive the average speed of these vehicles. To achieve this, we utilize a straightforward curve-fitting technique known as polynomial fitting, defined as follows:
$y(x,\omega)=\omega_{0}+\omega_{1}x+\omega_{2}x+\ldots+w_{M}x^{M}=\sum_{j=0}^{M}\omega_{j}x^{j}$ (11)
where m is the order of the polynomial, $\omega=(\omega_{0},...\omega_{M})^{T}$ represents the polynomial coefficients as a vector. Our objective is to obtain the coefficients of the fitting polynomial by minimizing the error between the fitting function and the data points in the training set: observations of $(x,t)$, $x=(x_{0},...x_{N})^{T}$ and $t=(t_{0},...t_{N})^{T}$. We employ the mean square error function, expressed as:
$E(\omega)=\frac{1}{2}\sum_{n=1}^{N}(y(x_{n},\omega)-t_{n})^{2}$ (12)
where the error function is quadratic with respect to the coefficient $\omega$ and has a non-negative value. When the error function is minimized, the unique solution $w^{*}$ along with the fitting polynomial $y(x,w^{*})$ can be obtained.
The vehicle trajectory extraction algorithm flow is shown in Algorithm 2.
Input: $A=(a_{ij})$: original optical fiber matrix; $i=1,...,m$, $j=1,...,n$, $a_{ij}\in\mathbb{R^{+}}$: selected time and distance.
Output: $\Sigma^{(S)}=\{(\sigma^{(s)})_{1\leq s\leq S}\}$: vehicle trajectory point sets.
1 Apply Algorithm 1 to $A=(a_{ij})$, returning preprocessed optical fiber matrix $D=(d_{ij})$; vehicle counts S and appearance time of each vehicle $k_{s}$. $(k_{s}+\alpha,l_{k_{s}+\alpha})$ is a pair of time and distance position points of each extracted vehicle trajectory;
2 $l_{k_{s}+\alpha}\in\mathbb{N^{+}}$ is distance position point corresponding to time position point $k_{s}+\alpha$ in $D=(d_{ij})$;
3 $\Sigma^{(S)}=\{(\sigma^{(s)})_{1\leq s\leq S}:\sigma^{(s)}=(k_{s}+\alpha,l_{k_{s}+\alpha}),\alpha\in\mathbb{N}$};
4 Select the initial velocity interval $v_{\min}^{(1)}$ to $v_{\max}^{(1)}$;
5 Calculate the initial distance interval $x_{\min}^{(1)}$ to $x_{\max}^{(1)}$;
6 foreach s = 1 to S; $\sigma^{(s)}\in\Sigma^{(S)}$ do
7       Add the first pair $(k_{s},l_{k_{s}})$ to vehicle trajectory point sets $\sigma^{(s)}$, where $l_{k_{s}}=1$;
8       $l_{k_{s}+1}$leftarrow ColIndex(Max ($D_{k_{s}+1,l_{k_{s}}+x_{\min}^{(1)}:l_{k_{s}}+x_{\max}^{(1)}}$));
9       if $k_{s}+1<m$ $\rm and$ $l_{k_{s}+1}<n$ then
10             Add pair $(k_{s}+1,l_{k_{s}+1})$ to each vehicle trajectory point set $\sigma^{(s)}$;
11
12      else
13             break;
14
15foreach s = 1 to S; $\sigma^{(s)}\in\Sigma^{(S)}$ do
16       for $\alpha\in\mathbb{N^{+}}$ do
17             $y^{(s)}$ = PolynomialFitting($\sigma^{(s)}$);
18             $v^{(s)}$ = Slope($y^{(s)}$);
19             Select the right level of confidence $\rm cof$;
20             Calculate velocity interval $(1+\rm cof)$$v^{(s)}$ to $(1+\rm cof)$$v^{(s)}$;
21             Calculate distance interval $x_{\min}^{(s)}$ to $x_{\max}^{(s)}$;
22             $l_{k_{s}+\alpha}$leftarrow ColIndex(Max ($D_{k_{s}+\alpha,l_{k_{s}+\alpha-1}+x_{\min}^{(s)}:l_{k_{s}+\alpha-1}+x_{\max}^{(s)}}$));
23             if $k_{s}+\alpha<m$ $\rm and$ $l_{k_{s}+\alpha}<n$ then
24                   Add pair $(k_{s}+\alpha,l_{k_{s}+\alpha})$ to each vehicle trajectory point set $\sigma^{(s)}$;
25
26            else
27                   break;
28
29
Return: Vehicle trajectory point sets $\Sigma^{(S)}$.
Algorithm 2 Trajectory Extraction Algorithm.
Based on the extraction results of key points for a single vehicle, we can deduce the vehicle’s approximate uniform-speed trajectory through polynomial fitting, as illustrated in Fig. 7.
The workflow diagram is shown in Fig. 5.
Figure 7: Trajectory extraction results. (a) trajectory key points of a large truck, and (b) trajectory fitting results of it.
IV Experiments
IV-A Datasets
In this study, we conducted data acquisition and analysis in two real highway scenarios, with a particular focus on a tunnel scenario. As our optical fiber is laid on only one side of the road, the target areas for this paper are explicitly stated. The algorithm proposed aims to extract vehicle trajectories within these target areas.
Target area in Jingjintang expressway dataset: Located in Junliang City Test Site, Dongli District, Tianjin, China, Jingjintang Expressway is a dual carriageway with three lanes in each direction. Our target area encompasses three lanes with a shoulder on the right, excluding the opposite lanes.
Target area in Jiurui tunnel dataset: The Jiujiang-Ruichang Expressway, situated in Jiujiang City, Jiangxi Province, China, represents a real highway tunnel scenario with a single carriageway featuring two lanes. Our target area focuses on the right lane, disregarding the left lane.
To assess the effectiveness of the optical fiber equipment, cameras were installed on the same right side of the road for data comparison. It is important to note that camera data is used here solely to evaluate the reliability of data collected by optical fiber. The algorithm proposed in this paper exclusively relies on optical fiber data and is independent of camera data.
IV-A1 Jingjintang Expressway Dataset
We selected a 320-meter section of Jingjintang Expressway for our study. Simultaneously, a camera mounted on the road captures movement information of the vehicles. Positioned at the starting point of the optical fiber laying side, the camera primarily monitors vehicles in the target area, as depicted in Fig. 8.
Figure 8: Camera detection of the target area, including 3 lanes and a shoulder.
When a vehicle traverses the area, the Distributed Acoustic Sensing (DAS) system records the corresponding vibration signals. To present the vehicles’ trajectory more intuitively, we utilize waterfall diagrams to plot the collected signal values, as illustrated in Fig. 9.
Figure 9: Optical fiber waterfall diagrams of vehicles passing at intervals in Jingjintang expressway dataset. (a) three trucks pass at intervals, (b) three cars pass at intervals, and (c) one truck and four cars pass at intervals.
In the waterfall diagram, the light blue color represents background noise, while prominently colored lines indicate vehicles. The fiber stripe extending from the top left to the bottom right corresponds to vehicles within the target area, and the stripe extending from the top right to the bottom left represents vehicles on the opposite side. Notably, the vehicle’s trajectory is depicted as an approximately straight line within the diagram. This is attributed to the fact that most vehicles on the highway maintain high and consistent velocities when traveling in lanes with unchanged speeds. In this representation, the intercept of the straight line signifies the starting position of the vehicle, and the absolute value of the slope represents its velocity.
While optical fiber waterfall diagrams excel at representing real-time vehicle location data, applying them in actual highway scenarios reveals certain inevitable challenges, as depicted in Fig. 10.
Figure 10: Displayed above are several representative waterfall diagrams featuring inaccurate optical fiber signals in the Jingjintang Expressway dataset. In the first row, optical fiber waterfall diagrams capture trucks passing nearly simultaneously in different lanes, significantly overlapping with neighboring cars. Moving to the second row, the diagrams illustrate vehicles from the opposite lane causing interference with the signals of vehicles in the target area. The presence of opposite vehicles is indicated by fiber stripes extending from the top right to the bottom left.
(1) Owing to high sensitivity and rapid response characteristics of optical fiber, large trucks are prone to generating large signal amplitudes. In instances where a car passes concurrently in close proximity, the car’s signal may become entirely submerged beneath the truck’s signal.
(2) Considering the relatively high number of large trucks on the highway, each truck generates substantial optical fiber signals. When multiple large vehicles traverse the roadway consecutively, they tend to interfere with each other, posing a challenge in distinguishing the count of vehicles within this cluster.
(3) While the signals from vehicles traveling in the opposite lane are not as clear as those within the target area, signals from trucks in the opposing lane can still interfere with the traffic signal data we aim to capture.
Additionally, we show a comparison between fiber stripes and camera ground truth to assess the accuracy of optical fiber. Fig. 11 presents the statistics of the number of trucks and cars for each minute within two hours. The full lines represent the ground truth detected by the camera, while the dashed lines represent the estimated counts detected by the optical fiber stripes. Comparing the results in Table. I, the optical fiber can capture nearly all the information about trucks. However, due to the aforementioned challenges leading to inaccurate optical fiber signals, the counting result for cars is lower than ground truth.
TABLE I: Comparison of real traffic flow and optical fiber stripe counts within 2 hours of Jingjintang expressway dataset.
Vehicle type Fiber stripes counts Camera counts Accuracy rate
Truck 241 254 94.88%
Car 381 734 51.90%
Total 622 988 62.95%
(a) Trucks.
(b) Cars.
Figure 11: Vehicle statistics counts in Jingjintang expressway dataset. The blue and orange lines show the counting results of the camera and the fiber stripe respectively.
IV-A2 Jiurui Tunnel Dataset
We also selected a test section in the Jiurui tunnel, including an 80-meter section and the entire 2000-meter tunnel.
Similarly, we draw optical fiber waterfall diagrams to get a more intuitive view of the vehicles in Fig. 12. Compared with Jingjintang expressway dataset, the optical fiber stripes in Jiurui tunnel dataset are more strong and obvious. This distinction arises from the tunnel’s configuration, consisting of just two lanes in the same direction, with trucks entering from the outer lane. Consequently, the distinct signals on the waterfall diagram correspond to trucks on the outer lane. Additionally, due to the attenuation effect of the tunnel, the signals from cars in the inner lane are too weak to be adequately recorded. Therefore, the relatively faint signals on the waterfall diagram represent cars in the inner lane.
Figure 12: Optical fiber waterfall diagrams of vehicles in Jiurui tunnel dataset of 80 meters section. (a) two trucks pass at intervals, (b) four trucks pass at intervals, and (c) two trucks and four cars pass at intervals.
The Jiurui tunnel dataset of short distance in Fig. 12 comprises 16 sets of original optical fiber data, with each set a collection time of 3-4 minutes and a collection distance of 80 meters. Among these, 7 datasets include corresponding video recordings. Comparing these optical fiber stripes in the optical fiber waterfall diagrams with the actual traffic flow in the videos in Fig. II, we observed that due to signal attenuation caused by the tunnel, the fiber stripes of small vehicles on the outer lane were faint, especially electric vehicles, which could not be adequately captured by the Distributed Acoustic Sensing (DAS) system. This explains the observed variance between optical fiber data acquisition and real traffic flow in the highway tunnel scenario.
TABLE II: Comparison of real traffic flow and optical fiber stripe counts in Jiurui tunnel dataset of short distance.
Collection time Fiber stripes counts Camera counts Accuracy rate
14:29-15:32 7 7 100.00%
15:40-15:43 6 6 100.00%
15:45-15:48 13 15 86.66%
16:01-16:04 6 7 85.71%
16:07-16:10 6 7 85.71%
16:19-16:22 5 5 100.00%
16:23-16:26 6 6 100.00%
In complex Jiurui tunnel long-distance scenarios shown in Fig. 2, prevalent conditions include low-speed limits and restrictions on lane changes and overtaking. Consequently, heavy traffic flows can lead to traffic congestion, as evident from the optical fiber waterfall diagrams. These diagrams reveal occurrences of multiple lines intersecting (congestion) and even overlapping (illegal overtaking) due to the specific traffic conditions and regulations enforced in the tunnel. The purpose of this paper is to track the whole trajectory of vehicles, even when the lane change problem is encountered.
IV-B Metrics
To make statistics on the characteristics of traffic flow, it is necessary to study the three fundamental indicators: velocity (V), flow (Q), and density (K) on road profile and road segment respectively.
On the road profile, traffic flow $\rm Q$ is defined as the number of vehicles N passing through a given profile of the road in a given period T: $\rm{Q=\frac{N}{T}}$. The average velocity of the vehicles passing through a given point is called Time Mean Speed (TMS). According to space-time accumulation, TMS on the profile should be the harmonic average velocity, which is a spatial concept. Therefore, TMS refers to the harmonic average of the location velocity of all vehicles passing through a profile of the road in a certain observation time: $\overline{\rm V}_{\rm{TMS}}=\frac{1}{\frac{1}{N}\sum_{i=1}^{N}\frac{1}{\rm V_{i}}}$. Density is expressed as occupancy rate, i.e., time density of vehicles, which represents the ratio of the time occupied by vehicles passing a certain road profile to the observation time in a certain observation time: $\rm K=\frac{\rm Q}{\overline{V}_{\rm{TMS}}}$.
On the road segment, density refers to the spatial density set of vehicles, that is, the number of vehicles N existing per unit road length L at a certain moment is $\rm K=\frac{\rm N}{\rm L}$. The average velocity of the vehicles in a given section is known as Space Mean Speed (SMS). According to space-time accumulation, SMS on the road segment should be the arithmetic average, which is a concept in time. Therefore, SMS refers to the arithmetic average of all vehicle instantaneous velocities in a certain observation section: $\overline{\rm V}_{\rm{SMS}}=\frac{\sum_{i=1}^{N}\rm V_{i}}{N}$. SMS is always lesser than or equal to TMS. Traffic flow refers to the ratio of the total distance of all vehicles to the accumulation of space and time in a certain period, $\rm Q=\rm K\overline{\rm V}_{\rm{SMS}}$.
IV-C Quantitative analysis
We tested the above algorithms on real highways. Our monitoring focuses on the target areas, and the algorithm proposed in this paper relies solely on optical fiber stripe data, independent of camera data. Consequently, when evaluating accuracy, our algorithm is compared only against optical fiber stripe data. By applying threshold processing, we can also distinguish between different vehicle classes, such as cars and trucks. Fig. 13 and Fig. 14 depict vehicle trajectory waterfall diagrams extracted by our algorithm with red lines indicating trucks and yellow lines indicating cars. Each line is uniquely identified with an individual ID.
Figure 13: An example of vehicle tracking results in the Jingjintang expressway dataset. (a) original waterfall diagram, (b) the red lines represent the extracted trajectory of trucks, and the yellow lines represent cars.
Figure 14: An example of vehicle tracking results in Jiurui tunnel dataset of short distance. (a) original waterfall diagram, (b) the red lines represent the extracted trajectory of trucks, and the yellow line represents a car.
Figure 15: Vehicle trajectory extraction results using the proposed method, Hough transform and Radon transform. The last two rows show the results using the Hough transform and Radon transform respectively.
Furthermore, We carried out comparative experiments with the Hough transform and Radon transform, employing the same preprocessing steps. As depicted in Fig. 15, our algorithm shows superior accuracy in extracting vehicle trajectories, even when these lines intersect or overlap, effectively distinguishing and separating individual vehicles.
We visually compared the algorithm results with optical fiber stripes in waterfall diagrams. The optical fiber dataset offers vehicle counts for both cars and trucks. Fig. 16 and Table. IV present a comparison of traffic flow data between the algorithm’s test results and the optical fiber results over two hours in the Jingjintang expressway dataset. Solid lines represent the vehicles detected by the algorithm, while dashed lines represent the optical fiber counts, which slightly deviate from the actual camera-based vehicle counts.
TABLE III: Comparison of algorithm test counts and optical fiber stripe counts in 7 sets of Jiurui tunnel dataset.
Collection time Algorithm test counts Fiber stripes counts Accuracy rate
14:29-15:32 5 7 71.42%
15:40-15:43 6 6 100.00%
15:45-15:48 11 13 84.61%
16:01-16:04 5 6 83.33%
16:07-16:10 5 6 83.33%
16:19-16:22 5 5 100.00%
16:23-16:26 6 6 100.00%
(a) Trucks
(b) Cars
Figure 16: Vehicle statistics counts in Jingjintang expressway dataset. The blue and orange lines show the counting results of the fiber stripe and the algorithm proposed in this paper respectively.
TABLE IV: Comparison of algorithm test counts and optical fiber stripe counts within 2 hours of Jingjintang expressway dataset.
Vehicle type Algorithm test counts Fiber stripe counts Accuracy rate
Truck 138 241 57.26%
Car 268 381 70.34%
Total 406 622 65.27%
The results from the Jiurui tunnel dataset are listed in Table. III. Among them, there were 143 instances of optical fiber stripes, and the algorithm’s detection results were 130 vehicles, yielding an accuracy of 90.90%. It is worth mentioning that the algorithm proposed in this paper not only exhibits consistency in counting but also maintains a high degree of alignment with vehicle trajectories through optical fiber stripes.
In addition to vehicle trajectory extraction, our approach can perform comprehensive statistical analyses, including velocity, flow and density on both road profiles and road segments. For example, based on the extraction results of structured vehicle data from the dataset in Fig. 14 of the tunnel highway scenario, we can specify a certain road profile or a certain segment and make traffic flow information statistics like Table. V.
TABLE V: Some traffic flow information statistics.
Time Position traffic flow information statistics
Vehicle counts Average velocity Flow Density
Road profile 0-1min 40m 4 56.38km/h 0.06 0.0011
1-2min 40m 3 59.93km/h 0.05 0.0008
2-3min 40m 2 59.27km/h 0.03 0.0005
3-4min 40m 6 54.07km/h 0.10 0.0018
Road segment 0-40m 30s 1 55.97km/h 0.55 0.0100
0-80m 30s 1 55.97km/h 0.27 0.0050
0-40m 150s 1 65.76km/h 0.65 0.0100
0-80m 150s 1 65.76km/h 0.32 0.0050
V Conclusion
In this paper, we have demonstrated the feasibility of employing optical fiber Distributed Acoustic Sensors(DAS). It can serve as a valuable supplement to existing traffic detection technologies and offers a feasible scheme for achieving long-distance vehicle trajectory tracking and comprehensive space-time traffic detection objectives for intelligent highways. However, there are certain deficiencies and areas for improvement in this research, including:
(1) Handing heavy traffic and parallel passages: Our proposed vehicle detection method works effectively for single-file traffic scenarios. However, in cases of heavy traffic with large and small vehicles passing in parallel, improving how we distinguish vehicle counts and precisely identify vibration signal positions is necessary.
(2) Dependency on vehicle starting points: The vehicle detection algorithm in this paper heavily relies on the starting points of vehicles. This approach may lead to missed detections when only using optical fiber data. In the era of multi-sensor fusion, integrating cameras could help determine the precise starting positions of vehicles, thereby improving trajectory accuracy.
(3) Considering complex traffic scenarios: To deploy fiber-assisted intelligent transportation successfully, we also need to address complex vehicle behaviors in real highway conditions, such as overtaking and stopping, to provide early warnings of emergencies.
Acknowledgments
This research was funded by Tianjin Yunhong Technology Development Grant No. 2021020531.
This work was supported in part by the National Natural Science Foundation of China Grant No. 61972404.